Question 7 Current Designs manufactures two different types of kayaks: rotomoulded kavaks and composte vayaks. The t folewing infarmation is aveilble for each product ine. Sales price/unit $1,200 $2,000 720 1,340 Variable costs/unit The companys fixed costs are s 1,001,040. An analysis of the sales mix identtes that retomolded kayaks-we ue 60%ofthe tot-ura Determine the weighted-average unit contribution margin for Current Designs Weighted-average unit contribution margin TO TEXT LENK TO TEXT LINK TO TEXT Determine the break-even point in units for Current Designs and idently how many units of each type of kwyak wil be told at 9e break-even pem (Ronan Break-even sales units Kayaks Composite Kayaks Break-even sales distribution units units LINK TO TEXT LINK TO TEXT LINK TO TEXY LINK TO TEXY LINK TO TET Assume that the sales mix changes, and rotonoulded kayaks now make up 70% of total units sold calculate the total number of units that would need to be sold to earn net income pn2 443,260 a d ident, hee many Required sales fb de iti-ach typ.. a-iii. units Kayaks Composite Kayaks unts units Break-even sales distributiorn
media%2F375%2F37539a6a-2076-4b67-bb2b-c3
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer   
Statementshowing Computations
Paticulars Rotomoulded Composite Total
Sales price per unit              1,200.00                 2,000.00
Less Variable Expenses per unit               (720.00)              (1,340.00)
Contribution Margin pr unit                  480.00                    660.00
% of sales 80% 20%
Weighted contribution                  384.00                    132.00                    516.00
Weighted contribution margin for current design is $516
b)
Fixed costs        1,001,040.00
Weighted contribution per unit                    516.00
Break even point in units =1001,040/516                 1,940.00
c)
Break even sales distribution =                  412.80                    103.20                    516.00
d)
Paticulars Rotomoulded Composite Total
Sales price per unit              1,200.00                 2,000.00
Less Variable Expenses per unit               (720.00)              (1,340.00)
Contribution Margin pr unit                  480.00                    660.00
% of sales 70% 30%
Weighted contribution                  336.00                    198.00                    534.00
Fixed costs        1,001,040.00
Desired profit        2,443,260.00
Desired contribution = 1001,040 + 2443,260        3,444,300.00
Required sales in units = 3444,300/534                 6,450.00
Sales distribution              4,515.00                 1,935.00
I have answered more than 4 parts as per HOMEWORKLIB RULES…. I could have answered last part as well but it is blurred. Not able to read it correctly